OpenStudy (anonymous):

Please help solve! 10e^(v+9)=12

OpenStudy (anonymous):

\[10e ^{v+9}=12\]

OpenStudy (isaiah.feynman):

Solving for v I presume.

OpenStudy (anonymous):

yes!

OpenStudy (isaiah.feynman):

Divide both sides by 10 so e is isolated.

OpenStudy (anonymous):

I got e^(v+9)=6/5

OpenStudy (anonymous):

is that correct?

OpenStudy (isaiah.feynman):

Yep. Now take the natural logarithm of both sides. So that v+9 comes down.

OpenStudy (anonymous):

would that be ln(v+9)=ln(6/5) ?

OpenStudy (anonymous):

no...@ali1029 ....remeber the formula....

OpenStudy (isaiah.feynman):

The right side is correct. There should be no "ln" on the left side

OpenStudy (anonymous):

so then it would just be: v+9 = ln(6/5) ?

OpenStudy (anonymous):

would that be similar to v+9=6/5 or does the ln on the right side have to remain there?

OpenStudy (isaiah.feynman):

What you want is